Joe earns $9.50 an hour, but he will get a 12% raise next week. What will be his hourly wage next week?
Answer by stanbon(75887)

New wage will be 9.5 + 0.12*9.5 = 1.12*9.5 = $10.64
==================
Cheers,
Stan H.